Determine the magnitude and direction of the force on an electron traveling 8.75x105 m/s horizontally to the east in a vertically upward magnetic field of strength 0.75 T.
Answer #1

Magnitude of force= qvB = 1.6*10-19*8.75*105*0.75 = 1.03*10-13 N

Direction would be south by right hand rule
Answer #2
Magnitude of force= qvB = 1.6*10-19*8.75*105*0.75 = 1.03*10-13 N
Direction would be south by right hand rule